# Why Do Open-Source LLMs Struggle with Data Analysis? A Systematic Empirical Study

### v2

@@ 4.3.4 Problem Diversity
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. The detailed methods and descriptions of the categories can be found in Appendix C. From the complete medium turn dataset, we select 2,220 examples under two settings: (1) a natural distribution that reflects the original domain frequencies, and (2) a balanced distribution that down-samples dominant domains while retaining all examples from underrepresented ones. Both settings preserve the overall difficulty distribution of the dataset. Results in Table 7 show minimal performance differences between the two settings, indicating that domain diversity alone does not significantly influence model performance in this context. These findings suggest that the effectiveness of a model is not solely determined by the diversity of problems it encounters during training. Rather, the diversity and richness of reasoning strategies—such as the depth of reasoning processes and the complexity of logical steps—appear to play a more influential role in shaping a model’s performance.
What Makes Data Effective ? The efficacy of large language models in data analysis is fundamentally shaped by the quality of the data rather than its sheer diversity. High-quality, well-structured tasks that emphasize nuanced complexity and transparent reasoning pathways are instrumental in refining the models’ analytical capabilities.
@@
Table 7: Analysis of Diversity characteristics using optimal turn settings (accuracy). The number of samples is 2220.
@@ 5 Towards Effective Data Analysis Performance
While the previous section examined how different aspects of strategic planning influence model performance, we now ask a practical question: how can these insights be leveraged to guide data collection and reuse in order to improve data analysis capabilities?
@@ 5.1 Strategy-Guided Data Synthesis
We design our data synthesis process in three systematically organized stages to construct a refined dataset that supports reasoning in data analysis. 1) Prompt-Based Answer Generation. We begin by leveraging prompt-based generation techniques to produce multiple candidate answers for each input query. This step ensures a diverse pool of responses, capturing a range of reasoning patterns and perspectives. 2) Targeted Instance Selection. Next, we prioritize medium-length dialogues and examples of medium to high difficulty, as these have been shown to facilitate more stable and effective learning. Through this filtering process, we select instances that strike a balance between complexity and informativeness. 3) Reasoning-Driven Data Enrichment. Finally, we enrich each selected instance with a concise reasoning summary. This step enhances abstraction and generalization by explicitly capturing the underlying reasoning process, allowing models to better learn transferable insights. By following this three-stage synthesis process with format standardization, we construct a dataset comprising 2.8k instances. The resulting dataset serves as the foundation for Supervised Fine-Tuning (SFT) to optimize model performance. A detailed overview of the synthesis process and the training parameters is provided in the Appendix D.
@@
Table 8: Comparison of Model Performances Across API and Different Sizes.
@@ 5.2 Evaluation Results
Table 8 presents the evaluation results of our approach, where we fine-tune 7B and 14B models on the curated dataset. Notably, the fine-tuned 7B model demonstrates substantial performance improvements compared to its baseline, while the fine-tuned 14B model achieves results that are comparable to or surpass those of GPT-4o. These results suggest that even simple, insight-driven adjustments to training data can yield substantial improvements in models’ performance on complex analytical tasks. However, the performance gains appear to diminish as model scale increases, suggesting a potential saturation point. One possible explanation is that our filtering strategy is constructed using the 7B model, making the resulting training distribution better aligned with its inductive biases. While this alignment benefits smaller architectures, it may be less effective for larger models with more diverse representational capacities. And a key limitation of our approach lies in the dataset itself. While the curated dataset provides value, it falls short in addressing the diversity and complexity required for more challenging tasks, making high-quality data a persistent bottleneck. To overcome this, future efforts should focus on expanding the dataset to include richer, more representative samples from real-world applications. Such improvements would not only better capture the variability and nuances of real-world scenarios but also enable further refinement of filtering strategies, enhancing scalability and generalizability across different model sizes and domains.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL). Prompt-based methods Zhao et al. (2024); Yao et al. (2023); Wang et al. (2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al., 2023; Zeng et al., 2024; Yin et al., 2024; Qiao et al., 2024; Song et al., 2024) provide task adaptation through labeled data, and have been applied in settings such as S1 Muennighoff et al. (2025), which introduces "budget forcing" to control reasoning steps, and LIMO Ye et al. (2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 DeepSeek-AI et al. (2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN Wang et al. (2025) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch Chen et al. (2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component. Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities Sun et al. (2024); Li et al. (2025). Data Copilot Zhang et al. (2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter Hong et al. (2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle Li et al. (2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench Hu et al. (2024), DSBench Jing et al. (2024), DA-code Huang et al. (2024), and DataSciBench Zhang et al. (2025) for assessing the effectiveness of LLM-based data analysis solutions.
@@ 7 Conclusion and Future Work
In this work, we conduct a detailed investigation into the data efficiency challenges faced by open-source LLMs in data analysis tasks. To this end, we collect a targeted dataset tailored to data analysis scenarios and explore the question: What strategies and resources are most effective for enhancing open-source LLMs on data analysis tasks? By examining model performance from the perspective of different capability dimensions, we identify key factors that influence effectiveness across varying interaction patterns. Our findings offer insights into how training strategies and multi-turn data design can be better aligned with the reasoning demands of data analysis tasks.
In the future, we plan to construct a large-scale synthetic dataset for data analysis and incorporate reinforcement learning to optimize data quality and further enhance model performance.
Limitations
While our experiments provide valuable insights into the performance of data efficiency in the context of data analysis, several limitations should be acknowledged. Our work does not incorporate experiments on a broader range of foundation models, such as LLaMA Dubey et al. (2024) and other emerging alternatives, which may exhibit distinct characteristics and behaviors. Second, due to GPU limitations, we were unable to extend our experiments to larger-scale datasets that could better reflect the diversity and complexity of real-world data analysis scenarios. Third, our evaluation is restricted to models of specific parameter sizes, leaving the behavior of significantly larger models (e.g., 72B) or smaller models(e.g., 2B) unexplored. Data analysis, as a domain, presents unique challenges compared to tasks like mathematical reasoning. These challenges include multi-turn interactions, sophisticated code reasoning, and external environment engagement, which demand data of higher diversity and complexity. Although we have not yet discovered domain-specific data capable of dramatically improving model performance—as has been the case in mathematical reasoning—our curated dataset provides valuable prior knowledge and new perspectives. Addressing these limitations in future work, through richer datasets and broader model testing, will be essential to advancing the scalability and adaptability of LLMs in data analysis and complex application domains.
@@
Task-Specific Details. We adopt different strategies based on the characteristics of each task. For Data Comprehension and Code Capability, we employ prompt-based evaluation methods as these tasks can be effectively assessed without additional training. In the case of Data Comprehension, we observe that when provided with file data, the model often redundantly attempts to reload the table using Python, leading to context overflow and reduced task accuracy. Moreover, fine-tuning may not reliably enhance the model’s ability to comprehend data and may inadvertently reinforce these redundant behaviors, making prompt-based evaluation a more suitable choice. For Code Capability, the current selection of models already supports effective evaluation of coding abilities, making fine-tuning unnecessary. For Strategic Planning, we use LoRA Hu et al. (2022) fine-tuning since this task requires the model to learn complex reasoning patterns and long-term dependencies through targeted training. To ensure fair comparisons across experiments, dataset sizes are controlled to be equal within the same task setting. However, some unavoidable differences exist across tasks due to varying data requirements and availability.
Training and Evaluation Details. For fine-tuning, we employ the LLaMA Factory framework (Zheng et al., 2024). Experiments are conducted on 4 NVIDIA A800 80G GPUs, leveraging the DeepSpeed ZeRO-3 optimization for efficient distributed training. We adopt a per-device batch size of 1, with gradient accumulation steps set to 4. Training is performed using bfloat16 precision, and a warmup ratio of 0.02. We employ a cosine learning rate scheduler with an initial learning rate of $1\times 10^{-5}$. All our fine-tuned models are evaluated using vLLM Kwon et al. (2023) with the temperature set to 0.
Appendix B Datasets Collection.
We synthesized a series of trajectories based on existing data. Notably, we exclude portions of DSBench data that involve image inputs to maintain consistency and ensure the focus of our analysis. Detailed statistics of the generated data can be found in Table 9.
@@
Appendix C Problem Diversity
@@
Appendix D Data Synthesis Process and Training Details.
@@
1:
  • Question set $\mathcal{Q}$: A collection of natural language queries to be processed.
  • File set $\mathcal{F}$: A set of files providing supplementary information.
2:Final responses $\mathcal{R}^{*}$: A structured set of high-quality responses, where each entry consists of a formatted response with embedded reasoning summary and its corresponding code segments.
3:Initialize an empty output set $\mathcal{R}^{*}\leftarrow\emptyset$
4:for each question $q\in\mathcal{Q}$ do
5:     $R_{q}\leftarrow\text{GenerateResponses}(q,\mathcal{F})$triangleright Generate initial candidate responses using question and supplementary files
6:     $R_{q}\leftarrow\text{FilterCorrectResponses}(R_{q})$
7:     if $R_{q}=\emptyset$ then
8:         continue
9:     end if
10:     for each response $r\in R_{q}$ do
11:         $r_{\text{clean}}\leftarrow\text{ApplyRulesAndFormat}(r)$triangleright Apply formatting rules to clean the response
12:         if $\text{IsLowComplexity}(r_{\text{clean}})$ or $\text{IsMediumTurn}(r_{\text{clean}})$ then
13:              continue
14:         end if
15:         $reason_{\text{chain}}\leftarrow\text{ExtractReasonChain}(r_{\text{clean}})$
16:         $reason_{\text{summary}}\leftarrow\text{Summarize}(reason_{\text{chain}})$
17:         $r_{\text{final}}\leftarrow\text{InsertBeforeCode}(r_{\text{clean}},reason_{\text{summary}})$
18:         Add $r_{\text{final}}$ to $\mathcal{R}^{*}$
19:     end for
20:end for
21:return $\mathcal{R}^{*}$triangleright Return the set of finalized responses

### v3

@@ 4.3 Strategic Planning
To quantify diversity, we employed a three-stage procedure to classify each question into distinct domains. The detailed methods and descriptions of the categories can be found in Appendix B. From the complete medium turn dataset, we select 2,220 examples under two settings: (1) a natural distribution that reflects the original domain frequencies, and (2) a balanced distribution that down-samples dominant domains while retaining all examples from underrepresented ones. Both settings preserve the overall difficulty distribution of the dataset. Results in Table 7 show minimal performance differences between the two settings, indicating that domain diversity alone does not significantly influence model performance in this context. These findings suggest that the effectiveness of a model is not solely determined by the diversity of problems it encounters during training. Rather, the diversity and richness of reasoning strategies—such as the depth of reasoning processes and the complexity of logical steps—appear to play a more influential role in shaping a model’s performance.
What Makes Data Effective ?
The efficacy of large language models in data analysis is fundamentally shaped by the quality of the data rather than its sheer diversity. High-quality, well-structured tasks that emphasize nuanced complexity and transparent reasoning pathways are instrumental in refining the models’ analytical capabilities.
@@
Table 7: Performance under different sampling strategies (% accuracy).
@@ 5 Towards Effective Data Analysis Performance
While the previous section examined how different aspects of data analysis influence model performance, we now ask a practical question: how can these insights be leveraged to guide data collection and reuse in order to improve data analysis capabilities?
@@ 5.1 Strategy-Guided Data Synthesis
We design our data synthesis process in three systematically organized stages to construct a refined dataset that supports reasoning in data analysis. 1) Prompt-Based Answer Generation. We begin by leveraging prompt-based generation techniques to produce multiple candidate answers for each input query. This step ensures a diverse pool of responses, capturing a range of reasoning patterns and perspectives. 2) Targeted Instance Selection. Next, we prioritize medium-length dialogues and examples of medium to high difficulty, as these have been shown to facilitate more stable and effective learning. Through this filtering process, we select instances that strike a balance between complexity and informativeness. 3) Reasoning-Driven Data Enrichment. Finally, we enrich each selected instance with a concise reasoning summary. This step enhances abstraction and generalization by explicitly capturing the underlying reasoning process, allowing models to better learn transferable insights. By following this three-stage synthesis process with format standardization, we construct a dataset comprising 2.8k instances. The resulting dataset serves as the foundation for Supervised Fine-Tuning (SFT) to optimize model performance. A detailed overview of the synthesis process and the training parameters is provided in Appendix C.
@@
Table 8: Performance comparison of our method and baselines across model types and sizes (% accuracy).
@@ 5.2 Evaluation Results
Table 8 presents the evaluation results of our approach, where we fine-tune 7B and 14B models on the curated dataset. Notably, the fine-tuned 7B model demonstrates substantial performance improvements compared to its baseline, while the fine-tuned 14B model achieves results that are comparable to or surpass those of GPT-4o. These results suggest that even simple, insight-driven adjustments to training data can yield substantial improvements in models’ performance on complex analytical tasks. However, the performance gains appear to diminish as model scale increases, suggesting a potential saturation point. One possible explanation is that our strategy is constructed using Qwen2.5-7B, making the resulting training distribution better aligned with its inductive biases. While this alignment benefits smaller architectures, it may be less effective for larger models with more diverse representational capacities. And a key limitation of our approach lies in the dataset itself. While the curated dataset provides value, it falls short in addressing the diversity and complexity required for more challenging tasks, making high-quality data a persistent bottleneck. To overcome this, future efforts should focus on expanding the dataset to include richer, more representative samples from real-world applications. Such improvements would not only better capture the variability and nuances of real-world scenarios but also enable further refinement of filtering strategies, enhancing scalability and generalizability across different model sizes and domains.
@@ 6 Related Work
To adapt LLMs to complex reasoning tasks, recent work has explored three main approaches: prompt engineering, supervised fine-tuning (SFT), and reinforcement learning (RL). Prompt-based methods (Zhao et al. 2024; Yao et al. 2023; Wang et al. 2023) improve reasoning performance by reformulating inputs to better elicit the model’s latent capabilities. SFT-based methods (Chen et al. 2023; Zeng et al. 2024; Yin et al. 2024; Qiao et al. 2024; Song et al. 2024) provide task adaptation through labeled data, and have been applied in settings such as S1 (Muennighoff et al. 2025), which introduces "budget forcing" to control reasoning steps, and LIMO (Ye et al. 2025), which demonstrates that complex mathematical abilities can emerge from carefully curated examples. RL-based approaches further align model behavior with human preferences and problem-solving strategies. Reinforcement learning further enhances reasoning capabilities by optimizing multi-step decision-making. For example, DeepSeek-R1 (DeepSeek-AI et al. 2025) leverages Group Relative Policy Optimization (GRPO) to train LLMs on complex reasoning tasks. RAGEN (Wang et al. 2025) introduces a modular RL framework that supports stable multi-turn learning via the StarPo architecture. In addition, ReSearch (Chen et al. 2025) integrates RL with retrieval-augmented generation (RAG), treating search operations as part of the reasoning process and learning when and how to retrieve external information.
@@
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component. Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities (Sun et al. 2024; Li et al. 2025). Data Copilot (Zhang et al. 2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter (Hong et al. 2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle (Li et al. 2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench (Hu et al. 2024), DSBench (Jing et al. 2024), DA-code (Huang et al. 2024), and DataSciBench (Zhang et al. 2025) for assessing the effectiveness of LLM-based data analysis solutions.
@@ 7 Conclusion and Future Work
In this work, we conduct a detailed investigation into the data efficiency challenges faced by open-source LLMs in data analysis tasks. By curating a dataset specifically for data analysis scenarios, we systematically assess how data structure, interaction patterns, and training strategies influence model performance. Our findings highlight that careful multi-turn data design and appropriately structured training data are critical for enhancing LLM reasoning in data analysis. In future work, we will focus on expanding our dataset and exploring broader model evaluations to further advance LLMs for data analysis.
@@
Task-Specific Details. We adopt different strategies based on the characteristics of each task. For Data Comprehension and Code Capability, we employ prompt-based evaluation methods as these tasks can be effectively assessed without additional training. In the case of Data Comprehension, we observe that when provided with file data, the model often redundantly attempts to reload the table using Python, leading to context overflow and reduced task accuracy. Moreover, fine-tuning may not reliably enhance the model’s ability to comprehend data and may inadvertently reinforce these redundant behaviors, making prompt-based evaluation a more suitable choice. For Code Capability, the current selection of models already supports effective evaluation of coding abilities, making fine-tuning unnecessary. For Strategic Planning, we use LoRA (Hu et al. 2022) fine-tuning since this task requires the model to learn complex reasoning patterns and long-term dependencies through targeted training. To ensure fair comparisons across experiments, dataset sizes are controlled to be equal within the same task setting. However, some unavoidable differences exist across tasks due to varying data requirements and availability.
Datasets Collection. We synthesized a series of trajectories based on existing data. Notably, we exclude portions of DSBench data that involve image inputs to maintain consistency and ensure the focus of our analysis. Detailed statistics of the generated data can be found in Table 9.
Training and Evaluation Details. For fine-tuning, we employ the LLaMA Factory framework (Zheng et al. 2024). Experiments are conducted on 4 NVIDIA A800 80G GPUs, leveraging the DeepSpeed ZeRO-3 optimization for efficient distributed training. We adopt a per-device batch size of 1, with gradient accumulation steps set to 4. Training is performed using bfloat16 precision, and a warmup ratio of 0.02. We employ a cosine learning rate scheduler with an initial learning rate of $1\times 10^{-5}$. All our fine-tuned models are evaluated using vLLM (Kwon et al. 2023) with the temperature set to 0.
@@
Appendix B Problem Diversity
@@
Appendix C Data Synthesis Process and Training Details.
@@
Input: Question set $\mathcal{Q}$; File set $\mathcal{F}$
Output: Final responses $\mathcal{R}^{*}$ (each with formatted reasoning summary and code segments)
1: Initialize empty output set $\mathcal{R}^{*}\leftarrow\emptyset$
2:for each question $q\in\mathcal{Q}$ do
3:  $R_{q}\leftarrow\text{GenerateResponses}(q,\mathcal{F})$
4:  $R_{q}\leftarrow\text{FilterCorrectResponses}(R_{q})$
5:  if $R_{q}=\emptyset$ then
6:   continue
7:  end if
8:  for each response $r\in R_{q}$ do
9:   $r_{\text{clean}}\leftarrow\text{ApplyRulesAndFormat}(r)$
10:   if $\text{IsLowComplexity}(r_{\text{clean}})$ or $\text{IsMediumTurn}(r_{\text{clean}})$ then
11:    continue
12:   end if
13:   $reason_{\text{chain}}\leftarrow\text{ExtractReasonChain}(r_{\text{clean}})$
14:   $reason_{\text{summary}}\leftarrow\text{Summarize}(reason_{\text{chain}})$
15:   $r_{\text{final}}\leftarrow\text{InsertBeforeCode}(r_{\text{clean}},reason_{\text{summary}})$
16:   Add $r_{\text{final}}$ to $\mathcal{R}^{*}$
17:  end for
18:end for
19:return $\mathcal{R}^{*}$